### v1

# Asymptotic evaluation of the information processing capacity in reservoir computing

Abstract
The squared error normalized by the target output is known as the information processing capacity (IPC) and is used to evaluate the performance of reservoir computing (RC). Since RC aims to learn the relationship between input and output time series, we should evaluate the IPC for infinitely long data rather than the IPC for finite-length data. To evaluate the IPC for infinitely long data using the IPC for finite-length data, we use an asymptotic expansion of the IPC and the least-squares method. Then, we show the validity of our method by numerical simulations.
1 Introduction
Since many kinds of data, e.g. video, audio, and stock prices, have time correlation, machine learning of time-series data is an important issue. Recurrent neural networks (RNNs) can store past input by recursively connecting hidden nodes [1] and can approximate the relationship between input and output time series with arbitrary accuracy [2]. Backpropagation through time (BPTT) is mainly used to train RNNs, but it is difficult to optimize network parameters due to the gradient vanishing or the gradient explosion [3]. Many variants of RNNs, such as LSTM [4] and GRU [5], have been proposed to solve the difficulty of training and have been very successful. However, BPTT calculations become slower for longer training data.
An echo state network (ESN) [6] is a kind of RNNs, which can finish training quickly by fixing the recurrent connections at the initial value and optimizing only the linear transformation of the readout layer. Not limited to neural networks, a linear combination of nonlinear dynamical systems can be used to approximate the relationship between input and output time series and is called a reservoir computing (RC) system [7]. RC systems can also approximate the relationship between input and output time series with arbitrary accuracy [8, 9]. Since no optimization is performed other than the linear transformation, an RC system is often inferior in performance to LSTM and other methods. However, it has the advantage that training finishes quickly by only calculating the pseudoinverse matrix. Therefore, it is more convenient than LSTM and other methods in situations where the target to be optimized changes frequently.
The performance of an RC system is evaluated by the mean squared error or the squared error normalized by the target output, and the latter is called the information processing capacity (IPC) [10, 11]. The IPC ranges between 00 and $1$, representing the approximation accuracy. Since the goal of RC is to learn the relationship between input and output time series, it is necessary to evaluate the IPC for infinitely long data, not the IPC for finite-length data. The simplest method of estimating the IPC for infinitely long data is to calculate the IPC for sufficiently long data. However, their difference remains and should be removed as much as possible. In this paper, we estimate the IPC for infinitely long data using the asymptotic expansion of the IPC and the least-squares method.
This paper is organized as follows. Section 2 reviews RC and its performance index, the IPC. In Section 3, we derive the asymptotic form of the IPC, and in Section 4, we show the validity of our method by numerical simulations. Section 5 summarizes the paper.
2 Review of RC and the IPC
The dynamics of RC is expressed as follows.
$x_{t}=$ $f(x_{t-1},u_{t}),$ (1)
$y_{t}=$ $w^{\top}x_{t}+b,$ (2)
Here, $u_{t}\in\mathbb{R}^{d_{0}},x_{t}\in\mathbb{R}^{d_{1}},y_{t}\in\mathbb{R}^{d_{2}}$ are the values of the input, the hidden nodes, and the reservoir output at time $t$, respectively. To simplify the notation, we rewrite $(x_{t}^{\top},1)^{\top}$ as $x_{t}$, and $(w^{\top},b)^{\top}$ as $w$, and Eq. (2) is expressed by
$y_{t}=$ $w^{\top}x_{t}.$ (3)
To obtain $x_{t}$ from Eq. (1), the values of the hidden nodes at a certain time $-\tau\,(<t)$ are required as the initial value. The initial value is not optimized in RC. Instead, we employ $f$, which reduces the dependence on the initial value through the recursive equation Eq. (1), and take a sufficiently large $\tau$. Furthermore, in most cases, $f$ reduces the dependence on the past input. For example, the hidden node values of the ESN, a typical example of RC systems, are given by [12]
$x_{t}=$ $\tanh(v_{1}^{\top}x_{t-1}+v_{2}^{\top}u_{t-1}+c).$ (4)
The cost function for an RC system is usually given by the mean squared error between the reservoir output sequence $(y_{1},\ldots,y_{T})$ and the target output sequence $(\hat{y}_{1},\ldots,\hat{y}_{T})$ for the input sequence $(\ldots,u_{1},\ldots,u_{T})$,  [^1] Although not used in this paper, a regularization term for $w$ may be added to the cost function.
$\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}-y_{t}\|^{2}=\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}-w^{\top}x_{t}\|^{2}.$ (5)
The linear transformation of the readout layer, $w$, is optimized by
$w_{T}=\operatorname*{arg\,min}_{w}\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}-w^{\top}x_{t}\|^{2}.$ (6)
In addition to the mean squared error, a quantity called the IPC, defined below,
$1-\frac{\min_{w}\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}-w^{\top}x_{t}\|^{2}}{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}_{t}\|^{2}},$ (7)
is also used as a performance index for RC. Due to the normalization by the target output, the IPC ranges from 00 to $1$, representing the accuracy of the approximation. Since the purpose of RC is to learn the relationship between input and output time series from a finite-length input/output set (training data), the actual performance is given by Eq. (7) with $T\to\infty$. Therefore, we should estimate the IPC for infinitely long training data from the IPC for finite-length training data. The simplest estimation method is to use sufficiently long training data. However, this method has the problem that we cannot evaluate the deviation from the limit value.
3 Asymptotic expansion of the IPC
In this section, we estimate the IPC for infinitely long data using the asymptotic expansion. First, we summarize the RC dynamics by referring to Ref. [10, 8, 9]. From Eq. (1), the values of the hidden nodes of the RC system at time $t$ are determined by the initial value $x_{-\tau}$ and the input sequence $(u_{-\tau},\ldots,u_{t})$. As we have explained, the RC system is usually designed to reduce dependence on the initial value and the past input. Hence, taking $\tau\to\infty$, we can consider that $x_{t}$ is determined only by the input sequence $(u_{t-s})_{s=0}^{\infty}$. Due to time-independence of Eq. (1), if we give the same input sequence at two different times, $(u_{t-s})_{s=0}^{\infty}=(v_{t^{\prime}-s})_{s=0}^{\infty}$, the hidden node values at $t$ with the input sequence $(u_{t-s})_{s=0}^{\infty}$ and those at $t^{\prime}$ with $(v_{t^{\prime}-s})_{s=0}^{\infty}$ are the same. This means that the values of the hidden nodes are determined by a sequence of real numbers, and the hidden nodes of the RC system can be considered as the mapping from a real number sequence (input sequence) to a real vector (hidden node variables). To simplify the notation, in the following, the input sequence $(u_{t-s})_{s=0}^{\infty}$ is represented as $u^{\prime}_{t}$, and ${\cal U}\subseteq(\mathbb{R}^{d_{0}})^{\mathbb{N}}$ denotes the set of input sequences. We express the target output function as $\hat{y}\colon{\cal U}\to\mathbb{R}^{d_{2}}$ and the reservoir function as $x\colon{\cal U}\to\mathbb{R}^{d_{1}+1}$. We assume the components of $x$, namely, $x_{1},\ldots,x_{d},x_{d+1}$, are linearly independent.
Next, we consider the stochasticity of the input sequence. Following Ref. [10, 9], we assume that $U=(U_{t})_{t\in\mathbb{Z}}$ is a stationary ergodic process. We use the notation $U^{\prime}_{t}=(U_{s})_{s\leq t}$, and in particular $U^{\prime}=(U_{t})_{t\leq 0}$. The target output sequence is given by $\{\hat{y}(U^{\prime}_{1}),...,\hat{y}(U^{\prime}_{T}),...,\hat{y}(U^{\prime}_{T+T^{\prime}})\}$, and we divide it into training data $\{(U^{\prime}_{t},\hat{y}(U^{\prime}_{t}))\}_{t=1}^{T}$ and test data $\{(U^{\prime}_{t},\hat{y}(U^{\prime}_{t}))\}_{t=T+1}^{T+T^{\prime}}$. We assume $T^{\prime}=O(T)$ in this paper. The readout layer is optimized by the training data,
$w_{T}=$ $\min_{w}\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (8)
where we use the notation $\|x(U^{\prime}_{t})\|^{2}=x(U^{\prime}_{t})^{\top}x(U^{\prime}_{t})$. The IPCs for the training and the test data at $w_{T}$ is called the training IPC and the test IPC, respectively, which are given by
$C_{T}=$ $1-\frac{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})-w_{T}^{\top}x(U^{\prime}_{t})\|^{2}}{\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})\|^{2}}=1-\frac{l_{T}(w_{T})}{\mu_{T}},$ (9)
$C^{\prime}_{T,T^{\prime}}=$ $1-\frac{\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\|\hat{y}(U^{\prime}_{t})-w_{T}^{\top}x(U^{\prime}_{t})\|^{2}}{\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\|\hat{y}(U^{\prime}_{t})\|^{2}}=1-\frac{l^{\prime}_{T^{\prime}}(w_{T})}{\mu_{T^{\prime}}}.$ (10)
For later convenience, we express the numerators and denominators of Eqs. (9) and (10) as
$l_{T}(w)=$ $\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (11)
$l^{\prime}_{T^{\prime}}(w)=$ $\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (12)
$\mu_{T}=$ $\frac{1}{T}\sum_{t=1}^{T}\|\hat{y}(U^{\prime}_{t})\|^{2}.$ (13)
Now, we derive the asymptotic forms of the training and the test IPC. In this paper, the IPC for infinitely long training data is called the true IPC $C_{0}$, and $w$ at that time is called the true parameter $w_{0}$. Since RC and the optimized value $w_{T}$ can be considered a linear regression model and its optimal parameter, the asymptotic theory of linear regression models in Ref. [13] can be applied. First, we introduce the notation,
$l(w,U^{\prime}_{t})=$ $\|\hat{y}(U^{\prime}_{t})-w^{\top}x(U^{\prime}_{t})\|^{2},$ (14)
$\mu(U^{\prime}_{t})=$ $\|\hat{y}(U^{\prime}_{t})\|^{2},$ (15)
and assume
$\mathbb{E}[l(w,U^{\prime})]<\infty,\ \ \mathbb{E}[\mu(U^{\prime})]<\infty.$ (16)
Then, from the law of large numbers, $l_{T}(w)$ converges in probability to $l(w)=\mathbb{E}[l(w,U^{\prime})]$ for $T\to\infty$, and $\mu_{T}$ to $\mu_{0}=\mathbb{E}[\mu(U^{\prime})]$. We find that $C_{0}$ and $w_{0}$ can be expressed by
$C_{0}=$ $1-\frac{\min_{w}l(w)}{\mu_{0}}=1-\frac{l(w_{0})}{\mu_{0}},$ (17)
$w_{0}=$ $\operatorname*{arg\,min}_{w}l(w).$ (18)
Next, we introduce $\zeta(U^{\prime}_{t})=\mu(U^{\prime}_{t})-\mu_{0}$ and define the following quantities,
$I_{\infty}(w)=$ $\frac{1}{4}\sum_{t=0}^{\infty}{\rm Cov}(\nabla_{w}l(w,U^{\prime}_{0}),\nabla_{w}l(w,U^{\prime}_{t})^{\top})+\frac{1}{4}\sum_{t=1}^{\infty}{\rm Cov}(\nabla_{w}l(w,U^{\prime}_{0}),\nabla_{w}l(w,U^{\prime}_{-t})^{\top}),$ (19)
$J=$ $\frac{1}{2}\frac{\partial^{2}l(w)}{\partial w\partial w}\biggr{|}_{w=w_{0}},\ \ J_{ij,kl}=\mathbb{E}[x_{i}(U^{\prime})\,x_{k}(U^{\prime})\,\delta_{jl}],$ (20)
$V_{\zeta,\infty}=$ $\sum_{t=0}^{\infty}{\rm Cov}(\zeta(U^{\prime}_{0}),\zeta(U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(\zeta(U^{\prime}_{0}),\zeta(U^{\prime}_{-t})),$ (21)
$C_{l,\zeta}(w)=$ $\sum_{t=0}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0})\,,\zeta(U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0})\,,\zeta(U^{\prime}_{-t})),$ (22)
$V_{l,\infty}(w)=$ $\sum_{t=0}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0})\,,l(w,U^{\prime}_{t}))+\sum_{t=1}^{\infty}{\rm Cov}(l(w,U^{\prime}_{0})\,,l(w,U^{\prime}_{-t})).$ (23)
We further assume that $I_{\infty}(w),J,V_{\zeta,\infty},C_{l,\zeta}$(w), and $V_{l,\infty}(w)$ are finite matrices. Below, we express $I_{\infty}(w_{0}),C_{l,\zeta}(w_{0})$, and $V_{l,\infty}(w_{0})$ as $I_{\infty},C_{l,\zeta}$, and $V_{l,\infty}$, respectively. Then, the following quantities,
$\xi_{T}=$ $\frac{1}{2}J^{-1/2}\nabla_{w}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}(l(w,U^{\prime}_{t})-l(w))\Bigr{|}_{w=w_{0}}=\frac{1}{2\sqrt{T}}J^{-1/2}\sum_{t=1}^{T}\nabla_{w}l(w_{0},U^{\prime}_{t})\,,$ (24)
$\xi^{\prime}_{T^{\prime}}=$ $\frac{1}{2}J^{-1/2}\nabla_{w}\frac{1}{\sqrt{T^{\prime}}}\sum_{t=T+1}^{T+T^{\prime}}(l(w,U^{\prime}_{t})-l(w))\Bigr{|}_{w=w_{0}}=\frac{1}{2\sqrt{T^{\prime}}}J^{-1/2}\sum_{t=T+1}^{T+T^{\prime}}\nabla_{w}l(w_{0},U^{\prime}_{t})\,,$ (25)
converge in distribution,
[…]
Following Ref. [14], the asymptotic forms can be derived from the mean-value theorem. From the mean-value theorem, there exists $c\in[0,1]$ which satisfies
$0=$ $\nabla_{w}l_{T}(w_{T})=\nabla_{w}l_{T}(w_{0})+\nabla_{w}^{2}l_{T}(w^{\prime})\,(w_{T}-w_{0}),$ (28)
$w^{\prime}=$ $cw_{T}+(1-c)w_{0}.$ (29)
Substituting $\nabla_{w}l_{T}(w_{0})=2(J/T)^{1/2}\xi_{T}$ and $\nabla_{w}^{2}l_{T}(w^{\prime})=J+o_{p}(1)$ into Eq. (28), we obtain the asymptotic form of $w_{T}$,
$w_{T}=w_{0}-(TJ)^{-1/2}\,\xi_{T}+o_{p}\left(\frac{1}{\sqrt{T}}\right).$ (30)
The third and higher order derivatives of $l_{T}$ is 00. Hence, applying the mean-value theorem to Eqs. (11), we find that
$l_{T}(w_{0})=$ $l_{T}(w_{T})+\nabla_{w}l_{T}(w_{T})^{\top}(w_{0}-w_{T})+\frac{1}{2}(w_{0}-w_{T})^{\top}\,\nabla_{w}^{2}l_{T}(w_{T})\,(w_{0}-w_{T}).$ (31)
Substituting $\nabla_{w}l_{T}(w_{T})=0$, $\nabla_{w}^{2}l_{T}(w_{T})=2J$ and Eq. (30) into Eq. (31), we obtain
$l_{T}(w_{T})=l_{T}(w_{0})-\frac{1}{T}\|\xi_{T}\|^{2}+o_{p}\left(\frac{1}{T}\right).$ (32)
Similarly, applying the mean-value theorem to Eq. (12), we find that
$l^{\prime}_{T^{\prime}}(w_{T})=$ $l^{\prime}_{T^{\prime}}(w_{0})+\nabla_{w}l^{\prime}_{T^{\prime}}(w_{0})^{\top}(w_{T}-w_{0})+\frac{1}{2}(w_{T}-w_{0})^{\top}\,\nabla_{w}^{2}l^{\prime}_{T^{\prime}}(w_{0})\,(w_{T}-w_{0}).$ (33)
Substituting $\nabla_{w}l^{\prime}_{T^{\prime}}(w_{0})=2(J/T^{\prime})^{1/2}\xi^{\prime}_{T^{\prime}}+o_{p}(1/\sqrt{T^{\prime}})$, $\nabla_{w}^{2}l^{\prime}_{T^{\prime}}(w_{0})=2J$ and Eq. (30) into Eq. (33), we obtain
$l^{\prime}_{T^{\prime}}(w_{T})=$ $l^{\prime}_{T^{\prime}}(w_{0})+\frac{1}{T}\|\xi_{T}\|^{2}+\frac{2}{\sqrt{TT^{\prime}}}\,\xi_{T}^{\top}\,\xi^{\prime}_{T^{\prime}}+o_{p}\left(\frac{1}{T}\right)\,.$ (34)
From the central limit theorem [15], the denominators of Eqs. (9) and (10) converge in distribution,
[…]
Substituting Eqs. (32), (34) and (35) into Eqs. (9) and (10), we obtain the asymptotic forms of the training and the test IPC,
$C_{T}=$ $1-\frac{1}{\mu_{0}}\left[l_{T}(w_{0})-\frac{l_{T}(w_{0})\zeta_{T}}{\mu_{0}}+\frac{l_{T}(w_{0})\zeta_{T}^{2}}{\mu_{0}^{2}}-\frac{\|\xi_{T}\|^{2}}{T}\right]+o_{p}\left(\frac{1}{T}\right),$ (36)
$C^{\prime}_{T,T^{\prime}}=$ $1-\frac{1}{\mu_{0}}\left[l^{\prime}_{T^{\prime}}(w_{0})-\frac{l^{\prime}_{T^{\prime}}(w_{0})\zeta_{T^{\prime}}}{\mu_{0}}+\frac{l^{\prime}_{T^{\prime}}(w_{0})\zeta_{T^{\prime}}^{2}}{\mu_{0}^{2}}+\frac{\|\xi_{T}\|^{2}}{T}+\frac{2\xi_{T}^{\top}\xi^{\prime}_{T^{\prime}}}{\sqrt{TT^{\prime}}}\right]+o_{p}\left(\frac{1}{T}\right).$ (37)
Taking expectations for the training and the test data, the mean and the variance of the training IPC are given by
$\mathbb{E}[C_{T}]=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{C_{l,\zeta}}{\mu_{0}^{2}}-\frac{l(w_{0})\,V_{\zeta,\infty}}{\ mu_{0}^{3}}+\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+o\left(\frac{1}{T}\right)\,,$ (38)
$\mathbb{V}[C_{T}]=$ $\frac{1}{T}\left[\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{l(w_{0})^{2}\,V_{\zeta,\infty}}{\mu_{0}^{4}}-\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\zeta}\right]+o\left(\frac{1}{T}\right)\,.$ (39)
Similarly, the mean and the variance of the test IPC are given by
$\mathbb{E}[C^{\prime}_{T,T^{\prime}}]=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{C_{l,\zeta}}{\mu_{0}^{2}}-\frac{T}{T^{\prime}}\,\frac{l(w_{0})\,V_{\zeta,\infty}}{\mu_{0}^{3}}-\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+o\left(\frac{1}{T}\right)\,,$ (40)
$\mathbb{V}[C^{\prime}_{T,T^{\prime}}]=$ $\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{T}{T^{\prime}}\,\frac{l(w_{0})^{2}\,V_{\zeta,\infty}}{\mu_{0}^{4}}-\frac{T}{T^{\prime}}\,\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\zeta}\right]+o\left(\frac{1}{T}\right)\,.$ (41)
Note that we have assumed $T^{\prime}=O(T)$. From Eqs. (38) and (40), we find that both means approach the true IPC, $C_{0}=1-l(w_{0})/\mu_{0}$, in $O(1/T)$. To estimate the true IPC, we first approximate the expectation values, $\mathbb{E}[C_{T}]$ and $\mathbb{E}[C^{\prime}_{T,T^{\prime}}]$, by the sample mean at each $T$. Then we ignore the term $o(1/T)$ in Eqs. (38) and (40) and use the least squares method to estimate the true IPC (see Appendix). Our method can be applied to RC systems that satisfy the assumption that $U$ is a stationary ergodic process, $\mathbb{V}[l(w,U^{\prime})]<\infty$, and $I_{\infty}(w),J,V_{\zeta,\infty},C_{l,\zeta}(w)$, and $V_{l,\infty}(w)$ are finite matrices. In the next section, we apply this method to several models.
Finally, we note that fitting the variances by Eqs. (39) and (41) fails when the true IPC is 00. In this case, the true parameter is $w_{0}=0$, and we find $l(w_{0},U^{\prime})=\|\hat{y}(U^{\prime})\|^{2}$. Thus, we obtain
${\rm Cov}(l(w_{0},U^{\prime}_{0}),l(w_{0},U^{\prime}_{t}))={\rm Cov}(l(w_{0},U^{\prime}_{0}),\|\hat{y}(U^{\prime}_{t})\|^{2})={\rm Cov}(l(w_{0},U^{\prime}_{0}),\zeta(U^{\prime}_{t}))={\rm Cov}(\zeta(U^{\prime}_{0}),\zeta(U^{\prime}_{t})).$ (42)
That is, $V_{l,\infty}=V_{\zeta,\infty}=C_{l,\zeta}$. Since the coefficients of the $1/T$ terms in Eqs. (39) and (41) vanish, fitting the coefficients of the $1/T$ terms fails. In the next section, we will show an example of the failure.
4 Numerical simulation
First, we show the effectiveness of our method in a system where each term of the asymptotic forms except for $o(1/T)$ can be precisely calculated. Next, we estimate the true IPCs in systems where the target outputs are given by Legendre polynomials. Finally, we apply our method to the estimation of the true IPC for the NARMA10 task.
4.1 A simple model
In this model, $U_{t}$ is a random variable from ${\rm Uniform}(-1,1)$. The hidden node of the RC system and the target output are given by
$x(U^{\prime})=\sum_{s=0}^{\infty}2^{-s}\,U_{-s},\ \ \hat{y}(U^{\prime})=1+x(U^{\prime}).$ (43)
The mean squared error is
$l(w)=\mathbb{E}[(\hat{y}(U^{\prime})-w\,x(U^{\prime}))^{2}]=\frac{4}{9}\,(w-1)^{2}+1.$ (44)
Thus, the true parameter is $w_{0}=1$, and the minimum value of the mean squared error is $l(w_{0})=1$. The true ICP is
$C_{0}=1-\frac{l(w_{0})}{\mu_{0}}=\frac{4}{13}.$ (45)
The other terms in the asymptotic expansions are
$C_{l,\zeta}=0,\ \ V_{l,\infty}=0,\ \ V_{\zeta,\infty}=\frac{6992}{1215},\ \ I_{\infty}=\frac{4}{3},\ \ J=\frac{4}{9}.$ (46)
We performed a numerical simulation. Training and test data were generated $10000$ times for each data length $T$, and the training and the test IPC were calculated. Their means and variances were modeled as
$C(T)=a+\frac{b}{T},\ \ C^{\prime}(T,T^{\prime})=a+\frac{c}{T^{\prime}},\ \ V(T)=\frac{d}{T},\ \ V^{\prime}(T,T^{\prime})=\frac{d}{T^{\prime}},$ (47)
and $a,b,c,d$ were estimated using the least squares method. The exact and estimated values of the asymptotic expansion are shown in Table 1.
Table 1: The ground truth and the estimated values of the coefficients of the asymptotic forms are shown. We can find that the estimated values matched well to the ground truth.
a b c d
ground truth 4/13 1839/10985 66606/10985 188784/142805
estimation 0.3076476776827442 0.1685955855549023 5.9262712596285585 1.3418518658017338
We find that the least-squares estimation is successful. We plot the means and variances of the IPC in Fig. 1 and find that they are almost on the asymptote up to the term $1/T$. Next, we plotted on the log-log scale to verify the dominant $T$-dependence. To remove the constant term, we subtracted $C_{0}$ from the means of the IPCs. (After subtracting $C_{0}$, the 18th item in the mean of the training IPC was negative, and we removed it from the log-log scale.) Most of the samples are on the asymptote up to the term $1/T$. Although the means of the training IPC fluctuate more, this is likely due to a lack of samples.
(a) Mean of IPCs in the simple model
(b) Mean of IPCs after removal of the constant term on a log-log scale in the simple model
(c) Variance of IPCs in the simple model
(d) Variance of IPCs on a log-log scale in the simple model
Figure 1: The means and the variances of the training and the test IPC obtained by simulation in the simple system, and the theoretical lines obtained by the asymptotic expansions are plotted against $T$. Although the means of the training IPC fluctuate in the log-log plot, most samples are roughly on the theoretical lines, which indicates the effectiveness of the estimation.
4.2 Legendre polynomials
The RC system approximates a mapping from the input to output time series by the linear combination of nonlinear dynamics. To capture the set of the functions that RC can produce, it is useful to evaluate the IPCs of the orthogonal polynomials [10, 11]. In this subsection, the input sequence $U^{\prime}$ is sampled independently from ${\rm Uniform}(-1,1)$, and the target output is the product of Legendre polynomials,
$\hat{y}(U^{\prime}_{t})=\prod_{i=1}^{\infty}L_{s_{i}}(u_{t-i}),$ (48)
where, $L_{s_{i}}$ is the $s_{i}$-th Legendre polynomial, and the number of nonzero elements in $(s_{i})_{i=1}^{\infty}$ is finite.
We employed two target outputs. One was a short-term linear task $\hat{y}_{t}=L_{1}(u_{t-1})$, which is easy to approximate with the ESN, and the other was a long-term nonlinear task $\hat{y}_{t}=L_{15}(u_{t-5})$, which is difficult to approximate. The number of hidden nodes was $d_{1}=100$, and the ESN parameters, $v_{1},v_{2},c$, were randomly chosen from the normal distribution. The spectral radius of $v_{1}$ was set to $0.9$, and the proportion of nonzero elements was set to $0.7$. For each $T$, we generated data $1000$ times and calculated the training and the test IPC. Then, we estimated the asymptotic parameters. Figs. 2 and 3 show the results when the first- and 15th-order Legendre polynomials were used for the target outputs, respectively. The means and the variances of the true IPC for the first-order polynomial task are roughly on the asymptotic curve, indicating that the estimation was successful. In the 15th-order polynomial task, the asymptote fitted the mean IPC samples well, and the estimated true IPC was almost 00. The variances were significantly off the asymptote in the log-log graph, which reflects our mention in Section 3 that when the true IPC is 00, the variance approaches 00 faster than $1/T$.
(a) Mean of IPCs for the 1st order polynomial task
(b) Mean of IPCs after removal of the constant term on a log-log scale for the 1st order polynomial task
(c) Variance of IPCs for the 1st order polynomial task
(d) Variance of IPCs on a log-log scale for the 1st order polynomial task
Figure 2: The means and the variances of the IPC using uniformly distributed input and Legendre first-order output were plotted, along with the asymptote estimated from them. The simulation results were mostly on the theoretical line, indicating the effectiveness of the estimation.
(a) Mean of IPCs for the 15th order polynomial task
(b) Mean of IPCs after removal of the constant term on a log-log scale for the 15th order polynomial task
(c) Variance of IPCs for the 15th order polynomial task
(d) Variance of IPCs on a log-log scale for the 15th order polynomial task
Figure 3: The means and the variances of the IPC using uniformly distributed input and 15th-order Legendre output are plotted, along with the asymptote estimated from them. Although the means are roughly on the theoretical line, the variances are significantly off. Since the true IPC is nearly 00, we can find that the variance decays faster than $1/T$. (In this graph, it decays at approximately $1/T^{2}$.)
4.3 NARMA10
The NARMA10 task is widely used as a benchmark for RC [16]. The input sequence is sampled independently from the uniform distribution. The output sequence of the NARMA10 is given by
$\hat{y}_{t}=\alpha y_{t-1}+\beta y_{t-1}\sum_{i=1}^{10}y_{t-i}+\gamma u_{t}u_{t-9}+\delta.$ (49)
We used ${\rm Uniform}(0,0.2)$ as the input distribution and $(\alpha,\beta,\gamma,\delta)=(0.3,0.05,1.5,0.1)$ as NARMA10 parameters. The number of hidden nodes was $d_{1}=100$, and the ESN parameters $v_{1},v_{2},c$, were randomly chosen from the normal distribution. The spectral radius of $v_{1}$ was set to $0.9$, and the proportion of non-zero elements was set to $0.7$. For each $T$, data were generated $1000$ times, and the training and the test IPC were calculated. Then, we estimate the asymptotic parameters. The results are shown in Fig. 4. We can see that both the means and variances are almost on the asymptote.
(a) Mean of IPCs for the NARMA10 task
(b) Mean of IPCs after removal of the constant term on a log-log scale for the NARMA10 task
(c) Variance of IPCs for the NARMA10 task
(d) Variance of IPCs on a log-log scale for the NARMA10 task
Figure 4: The means and the variances of the IPC obtained from the NARMA10 task and the asymptote estimated from them are plotted. The simulation results are almost on the theoretical lines.
5 Summary
RNNs and their variants such as LSTM [4] and GRU [5] can approximate the relationship between the input and the output time series. However, the gradient learning takes a long time as data length becomes longer. RC is a system that approximates the relationship between input and output time series by the linear combination of nonlinear dynamics and finishes training quickly by optimizing only the linear combination.
The performance of RC is evaluated by the mean squared error or the squared error normalized by the target output. The latter is called the IPC [10, 11]. To know the approximation accuracy of the relationship between input and output time series, we should evaluate the IPC for infinitely long data, not the IPC for finite-length data. Simply evaluating the IPC with long data cannot remove the difference from the limit value. We propose a method to estimate the IPC for infinitely long data by using the asymptotic expansions of the training and the test IPC and the least-squares method. Then, we show the validity of our method by numerical simulations. Although we used ESNs in the demonstrations, our method can be applied to RC systems that satisfies the finiteness of the covariance matrices, as mentioned in the last second paragraph of Section 3.
Our method can estimate the IPC more accurately than simply extending the data length. However, it is computationally heavy due to multiple trials with long data. This is a trade-off with the estimation accuracy. A computationally less costly method for evaluating the IPC is our future work.
Acknowledgement
This paper is based on results obtained from a project, JPNP16007, commissioned by the New Energy and Industrial Technology Development Organization (NEDO). We appreciate Professor Takashi Morie and Doctor Tomoyuki Kubota for their fruitful comments.
Appendix A Least mean square
Let $f_{1}(n)=a+\frac{b_{1}}{n}$ and $f_{2}(n)=a-\frac{b_{2}}{n}$ be the asymptotic mean IPCs for the training and the test data, respectively, and $\{(n_{i,1},g_{i,1})\}_{i=1}^{N}$ and $\{(n_{i,2},g_{i,2})\}_{i=1}^{N}$ are the mean IPCs for the training and the test data at each data length. We determine $a,b_{1},b_{2}$ by the least mean square of the cost function below,
$L(a,b,b^{\prime})=\frac{1}{2}\sum_{i=1}^{N}\left[\left(a+\frac{b_{1}}{n_{i,1}}-g_{i,1}\right)^{2}+\left(a-\frac{b_{2}}{n_{i,2}}-g_{i,2}\right)^{2}\right].$ (50)
Using the notation,
$\alpha_{1}=\sum_{i=1}^{N}\frac{1}{n_{i,1}},\ \ \alpha_{2}=\sum_{i=1}^{N}\frac{1}{n_{i,2}},\ \ \beta_{1}=\sum_{i=1}^{N}\frac{1}{n_{i,1}^{2}},\ \ \beta_{2}=\sum_{i=1}^{N}\frac{1}{n_{i,2}^{2}},$ (51)
$s_{1}=\sum_{i=1}^{N}g_{i,1},\ \ s_{2}=\sum_{i=1}^{N}g_{i,2},\ \ t_{1}=\sum_{i=1}^{N}\frac{g_{i,1}}{n_{i,1}},\ \ t_{2}=\sum_{i=1}^{N}\frac{g_{i,2}}{n_{i,2}},$ (52)
we can obtain $a,b,b^{\prime}$ from
$\begin{pmatrix}2N&\alpha_{1}&-\alpha_{2}\\ \alpha_{1}&\beta_{1}&0\\ \alpha_{2}&0&-\beta_{2}\\ \end{pmatrix}\,\begin{pmatrix}a\\ b_{1}\\ b_{2}\\ \end{pmatrix}=\begin{pmatrix}s_{1}+s_{2}\\ t_{1}\\ t_{2}\\ \end{pmatrix}.$ (53)
In the same way, let $f^{\prime}_{1}(n)=\frac{d}{n}$ and $f^{\prime}_{2}(n)=\frac{d}{n}$ are the asymptotic IPC variances for the training and the test data, respectively, and $\{(n_{i,1},g^{\prime}_{i,1})\}_{i=1}^{N}$ and $\{(n_{i,2},g^{\prime}_{i,2})\}_{i=1}^{N}$ are the IPC variances at each data length. We determine $d$ by the least mean square,
$L^{\prime}(d)=\frac{1}{2}\sum_{i=1}^{N}\left[\left(\frac{d}{n_{i,1}}-g^{\prime}_{i,1}\right)^{2}+\left(\frac{d}{n_{i,2}}-g^{\prime}_{i,2}\right)^{2}\right],$ (54)
and obtain
$d=\frac{\sum_{i=1}^{N}\left(\frac{g^{\prime}_{i,1}}{n_{i,1}}+\frac{g^{\prime}_{i,2}}{n_{i,2}}\right)}{\sum_{i=1}^{N}\left(\frac{1}{n_{i,1}^{2}}+\frac{1}{n_{i,2}^{2}}\right)}=\frac{t^{\prime}_{1}+t^{\prime}_{2}}{\beta_{1}+\beta_{2}},$ (55)
where
$t^{\prime}_{1}=\sum_{i=1}^{N}\frac{g^{\prime}_{i,1}}{n_{i,1}},\ \ t^{\prime}_{2}=\sum_{i=1}^{N}\frac{g^{\prime}_{i,2}}{n_{i,2}}.$ (56)